OpenStudy (anonymous):

which expression is equivalent to ( -2)(a+6)?(1point) -2a+6 2a+12 -2a-12 -2a+12

OpenStudy (anonymous):

a(b+c) = ab+ ac

OpenStudy (anonymous):

ok?

OpenStudy (nincompoop):

not okay

OpenStudy (uri):

Multiply -2 with the whole bracket -2*a+ 2*6

OpenStudy (anonymous):

and you have a=-2 , b=a(not last a) , c=6

OpenStudy (anonymous):

and i think you can do yourself madam :)

OpenStudy (anonymous):

thanks uri

OpenStudy (nincompoop):